# Cognitive Limits Shape Language Statistics

### v1

@@ 2 Zipf’s Law and optimal compression
Figure 1: Statistical properties of languages. Top left: Normalized frequency-rank distribution $p(R)$ of single words (1-grams) for English, French, Italian, Spanish, and German. For each of these languages, the most frequent $\sim 8000$ words follow a power law with exponent -1, known as Zipf’s Law. In the tail of the distribution, the exponent shifts to 2 for most languages, indicating a transition from an optimal to a suboptimal regime. This shift reflects the need to form combinations of words to represent an increasingly complex conceptual space using a finite vocabulary. In German, however, the exponent bends less sharply, with a value of approximately 1.5. This reflects its strategy of expanding the vocabulary by creating new words from combinations of existing ones, as shown in the bottom panel. Top right: Frequency-rank distribution of n-grams (obtained aggregating from 1-grams to 5-grams) for the five languages. Here, n-grams are treated as representations of individual concepts, and the resulting distribution follows Zipf’s Law without deviation, indicating efficient concept representation through word combinations. Notably, similar words in related languages appear at comparable positions in the distribution. For instance, in German, the compound word “Wasserflasche” appears close in rank to the corresponding English bigram “water bottle”. This suggests that, despite differences in word formation strategies, these distributions directly reflect underlying conceptual representations. Plots are generated using n-gram frequencies from the Google Ngram database [46]. Bottom: In this panel, we illustrate the effect of using word combinations to represent complex concepts. Cognitive constraints limit the indefinite growth of vocabulary, leading to an increased reliance on word combinations as concepts grow more complex. This change impacts the word frequency distribution, causing a shift from an exponent of 1 to 2. German differs since it often creates compound words for complex concepts, thus retaining a lower exponent of $\approx$ 1.5. However, as words can be arbitrarily long as often happens in German, a larger vocabulary does not necessarily imply using fewer characters.
@@
The only way to cover an unlimited space of concepts with a finite vocabulary is to use combinations of words to represent increasingly complex meanings [54, 55]. This happens in all the natural languages, as illustrated in Fig. 1. Let’s consider each combination of one or more words as representing a unique concept. The frequency rank distribution of n-grams should follow Zipf’s Law, indicating an efficient representation of concepts [56, 57]. In Fig. 1, we show the frequency rank distribution of n-grams for different languages, which strictly follows Zipf’s Law in each case. These observations suggest a direct strategy to enhance language efficiency: identifying the most frequent n-grams and replacing combinations of more than one word with a single codeword. However, this would require a virtually infinite vocabulary, which is prevented by cognitive limitations in the human case. This argument also explains why German has not yet reached the regime exponent of $2$ typical of other languages, and exhibits a smaller exponent of $1.5$ in its tail. German achieves higher efficiency by associating unique words with complex concepts, often relying on explicit concatenation of many words.
@@ 3 Language network properties
Like many complex networks [66, 67], the WCN is largely scale-free, meaning that the degree distribution follows a power law [26, 33, 68]. A key feature of this network is that the nodes can be divided into two regions, each displaying different scaling exponents [69]. The “kernel” of the network, composed of the approximately 8000 most frequent words, follows $p(k)\sim k^{-\gamma}$ with $\gamma$ close to $3$, a value typical of Barabasi-Albert networks [35, 70]. This well-known structure offers great navigability, with a high clustering coefficient and a short average path length between nodes. Every word in the kernel can be reached from any other with fewer than three steps [36]. The preferential attachment mechanism ensures that frequent words are more accessible, a phenomenon known as the “frequency effect” [71]. This structure optimizes the retrieval of common words and is closely linked to the communication’s efficiency within the core lexicon.
@@ 4 A generative model for the evolution of language
We can measure $D_{0}$ by fitting the model to a real corpus of text, obtaining $D_{0}\approx 8000$, which aligns with previous estimates of the core lexicon size [26, 27]. To test the hypothesis that $D_{0}$ is related to cognitive capabilities, we compare texts produced by native English speakers (L1 corpus) with those written by non-native speakers learning the language (L2 corpus [84]). We show the results in Fig. 7 in the Appendix C. The L2 corpus exhibits a less efficient use of the language, reflected in a smaller size of the kernel lexicon. Having a smaller vocabulary, learners rely heavily on combinations of words to express complex ideas, producing longer texts with fixed information content. The value measured in the L2 corpus is $D_{0}\approx 200$, highlighting the reduced optimization capabilities of learners compared to native speakers.
@@ 5 Discussion
The idea that variations in Zipf’s exponent can differentiate writers with different cognitive or linguistic capabilities dates back to the first studies in linguistics [1] and has been extensively investigated [85]. Notably, individuals affected by schizophrenia tend to exhibit larger Zipf’s exponents. Our framework provides a direct explanation for these observations: these individuals show a disordered and less efficient use of language, which manifests as a smaller kernel lexicon, revealing the exponent $2$ much earlier than in typical speakers. In our framework, linguistic disorders and differences among speakers (such as the case of English learners reported in Section 4) can be detected by measuring the size of the kernel lexicon $D_{0}$. This approach allows for a direct comparison and differentiation of corpora produced by different speakers.
@@ Acknowledgments:
We acknowledge valuable discussions with M. Marsili, S. Thurner and G. De Marzo.
@@ C.2 Analytical solution of the model
C.3 The size of the kernel lexicon and comparison between L1 and L2 corpora
@@ Appendix D Data sources and other languages analysis
The frequency-rank distribution of words in Fig. 1 was retrieved from the version 20200117 of the Google Ngrams dataset [46]. We downloaded frequency data for 1- to 5-grams in English, Spanish, Italian, French, and German, filtering occurrences from 1950 onward. This dataset provides only word frequencies without access to the original text, though syntactic connections can be inferred from higher-order n-grams.
@@
The positions (rank) of the words highlighted in Fig 1, retrieved from the data, are: water, 243; Wasser, 268; bottle, 2546; Flasche, 3300; water bottle, 75241; Wasser Flasche, 76789.

### v2

@@ 2 Zipf’s Law and optimal compression
Figure 1: Statistical properties of languages. Top left: Normalized frequency-rank distribution $p(R)$ of single words (1-grams) for English, French, Italian, Spanish, and German. For each of these languages, the most frequent words follow a power law with exponent -1, known as Zipf’s Law. In the tail of the distribution, the exponent shifts to 2 for most languages, indicating a transition from an optimal to a suboptimal regime. This shift reflects the need to form combinations of words to represent an increasingly complex conceptual space using a finite vocabulary. In German, however, the exponent bends less sharply, with a value of approximately 1.5. This reflects its strategy of expanding the vocabulary by creating new words from combinations of existing ones [91], as shown in the bottom panel. Top right: Frequency-rank distribution of n-grams (obtained aggregating from 1-grams to 5-grams) for the five languages. Here, n-grams are treated as representations of individual concepts, and the resulting distribution follows Zipf’s Law without deviation, indicating efficient concept representation through word combinations. Notably, similar words in related languages appear at comparable positions in the distribution. For instance, in German, the compound word “Wasserflasche” appears close in rank to the corresponding English bigram “water bottle”. This suggests that, despite differences in word formation strategies, these distributions directly reflect underlying conceptual representations. Plots are generated using n-gram frequencies from the Google Ngram database [46]. Bottom: In this panel, we illustrate the effect of using word combinations to represent complex concepts. Cognitive constraints limit the indefinite growth of vocabulary, leading to an increased reliance on word combinations as concepts grow more complex. This change impacts the word frequency distribution, causing a shift from an exponent of 1 to 2. German differs since it often creates compound words for complex concepts, thus retaining a lower exponent of $\approx$ 1.5. However, as words can be arbitrarily long as often happens in German, a larger vocabulary does not necessarily imply using fewer characters.
@@
The only way to cover an unlimited space of concepts with a finite vocabulary is to use combinations of words to represent increasingly complex meanings [54, 55]. This happens in all the natural languages, as illustrated in Fig. 1. Let’s consider each combination of one or more words as representing a unique concept. The frequency rank distribution of n-grams should follow Zipf’s Law, indicating an efficient representation of concepts [56, 57]. In Fig. 1, we show the frequency rank distribution of n-grams for different languages, which strictly follows Zipf’s Law in each case. These observations suggest a direct strategy to enhance language efficiency: identifying the most frequent n-grams and replacing combinations of more than one word with a single codeword. However, this would require a virtually infinite vocabulary, which is prevented by cognitive limitations in the human case. This argument also explains why German has not yet reached the regime exponent of $2$ typical of other languages, and exhibits a smaller exponent of $1.5$ in its tail. German associates unique words with complex concepts, often relying on explicit concatenation of many words [91]. This reduces the need to rely on multi-word combinations, and thus slows down the exponent from shifting to $2$, keeping it around a smaller value.
@@ 3 Language network properties
Like many complex networks [66, 67], the WCN is largely scale-free, meaning that the degree distribution follows a power law [26, 33, 68]. A key feature of this network is that the nodes can be divided into two regions, each displaying different scaling exponents [69]. The “kernel” of the network, composed of the approximately 8000 most frequent words for many languages, follows $p(k)\sim k^{-\gamma}$ with $\gamma$ close to $3$, a value typical of Barabasi-Albert networks [35, 70]. This well-known structure offers great navigability, with a high clustering coefficient and a short average path length between nodes. Every word in the kernel can be reached from any other with fewer than three steps [36]. The preferential attachment mechanism ensures that frequent words are more accessible, a phenomenon known as the “frequency effect” [71]. This structure optimizes the retrieval of common words and is closely linked to the communication’s efficiency within the core lexicon.
@@ 4 A generative model for the evolution of language
We can measure $D_{0}$ by fitting the model to a real corpus of text, obtaining $D_{0}\approx 8000$, which aligns with previous estimates of the core lexicon size for the English language [26, 27]. To test the hypothesis that $D_{0}$ is related to cognitive capabilities, we compare texts produced by native English speakers (L1 corpus) with those written by non-native speakers learning the language (L2 corpus [84]). We show the results in Fig. 7 in the Appendix C. The L2 corpus exhibits a less efficient use of the language, reflected in a smaller size of the kernel lexicon. Having a smaller vocabulary, learners rely heavily on combinations of words to express complex ideas, producing longer texts with fixed information content. The value measured in the L2 corpus is $D_{0}\approx 200$, highlighting the reduced optimization capabilities of learners compared to native speakers. It is important to stress that this interpretation is meaningful only for intra-language comparisons. When comparing different languages, differences in $D_{0}$ must be interpreted with caution. For instance, German displays a significantly larger kernel lexicon ($D_{0}\approx 20000$) compared to other languages [27], but this does not indicate greater efficiency. It rather reflects the language’s strong tendency to form compound words [91], which artificially inflates the vocabulary size, as discussed in Appendix C.3.
@@ 5 Discussion
The idea that variations in Zipf’s exponent can differentiate writers with different cognitive or linguistic capabilities dates back to the first studies in linguistics [1] and has been extensively investigated [85]. Notably, individuals affected by schizophrenia tend to exhibit larger Zipf’s exponents. Our framework provides a direct explanation for these observations: these individuals show a disordered and less efficient use of language, which manifests as a smaller kernel lexicon, revealing the exponent $2$ much earlier than in typical speakers. In our framework, linguistic disorders and differences among speakers (such as the case of English learners reported in Section 4) can be detected by measuring the size of the kernel lexicon $D_{0}$. This approach allows for a direct comparison and differentiation of corpora produced by different speakers. However, cross-language comparisons must be treated with caution, as the size of the kernel lexicon can be influenced by structural properties of the language. In the case of German, for instance, the frequent use of compound words leads to a larger kernel lexicon without a corresponding gain in communicative efficiency.
@@ Acknowledgments:
We acknowledge valuable discussions with M. Marsili, S. Thurner, E. G. Altmann, and G. De Marzo.
@@ C.2 Analytical solution of the model
C.3 The size of the kernel lexicon, comparison between L1 and L2 corpora and cross-language comparison
@@
These considerations, however, require additional care when comparing across languages. Although in many languages the kernel lexicon of native speakers is around $D_{0}\sim 8000$, in some cases it appears significantly larger. For instance, German exhibits $D_{0}\sim 20000$. This would imply that German speakers are more than twice as “capable” as speakers of other languages, but such interpretation is misleading. As previously noted, German has a strong morphological tendency to form compound words [91]. Linguistic studies show that over 60% of the German lexicon consists of compounds [92], compared to only about 30% in English [93]. This morphological feature may account for the larger size of the German kernel lexicon, as also noted in [27].
If the larger size of $D_{0}$ is merely an artifact of compounding, it does not reflect a real gain in communicative efficiency. In practice, the additional German words included in the kernel lexicon are often concatenations of other base words which are equivalent to multi-word expressions in other languages, as illustrated in Fig. 1. German can thus express the same content using fewer word tokens [94]. However, because it does not introduce genuinely shorter representations of meaning, the total number of letters required remains approximately the same. The reduced word count is offset by longer average word length. Any saving in overall text length is therefore minimal.
By contrast, the L1/L2 comparison within the same language is conceptually different. Native speakers produce shorter texts in terms of word count compared to learners, due to their larger kernel lexicon $D_{0}$. In this case, a smaller number of words also results in a smaller number of letters, since both groups use the same lexicon and grammar. Therefore, a smaller $D_{0}$ directly reflects a reduced efficiency in language use.
@@ Appendix D Data sources and other languages analysis
The frequency-rank distribution of single words in Fig. 1 (top left) was retrieved from the version 20200117 of the Google Ngrams dataset [46]. The frequency-rank distribution of n-grams in Fig. 1 (top right) was instead retrieved, due to computational contraints, from version 20120701, which is smaller than 20200117. We downloaded frequency data for 1- to 5-grams in English, Spanish, Italian, French, and German, filtering occurrences from 1950 onward. The Google Ngrams provides only word frequencies without access to the original text, though syntactic connections can be inferred from higher-order n-grams.
The positions (rank) of the words highlighted in Fig 1, retrieved from the data, are: water, 318; Wasser, 305; bottle, 5312; Flasche, 6324; water bottle, 322242; Wasserflasche, 329067.